# Regulating Sommerfeld resonances for multi-state systems and higher partial waves

### v1

@@ 2.6.3 Finite range Coulomb potential
In both the $\ell=0$ and $\ell\geq 1$ cases, the divergent behavior of $C_{\ell}^{2}$ as $p\rightarrow 0$ is suppressed by the last factor on the right-hand side of this equation. In Fig. 9, we show the Sommerfeld enhancement as calculated by the conventional formula Eq. 45, the full formula Eq. 42, and the simplified formula Eqs. 99, 100. For the $s$-wave case, the behavior of the full formula is well captured by the simplified formula Eq. 99. For higher partial waves, the simplified formula Eq. 100 describes the behavior of the full formula fairly well, except for the points where $\tilde{z}_{\ell}$ crosses 00 and flips its sign; these behaviors are similar to those for the finite square well.
Figure 9: The ratio between the annihilation cross sections with and without the Sommerfeld effect for the finite-range Coulomb potential, for the $\ell=0$ (left), $\ell=1$ (middle) and $\ell=2$ (right) cases. The black solid curves shows the results computed using the conventional formula in Eq. 45, the red solid curves correspond to the full corrected result given in Eq. 42, and the blue dashed curves correspond to the simplified formula of Eqs. 99, 100. We take $k(p_{0})=(10^{3}+10^{3}i)a_{B}^{-1}$ for the $s$-wave case, $(10^{5}+10^{5}i)a_{B}^{-3}$ for the $p$-wave case, and $(10^{5}+10^{5}i)a_{B}^{-5}$ for the $d$-wave case. We choose $p=10^{-2}a_{B}^{-1}$ and $a=10^{-10}a_{B}$ in all cases.
@@ 3 The multi-state case for general $\ell$
Assuming the energy associated with the potential to be small compared to the mass, we will generally treat all states that can be excited due to the potential as having approximately equal masses $M$, and treat the mass differences between them as terms in the potential. In this case, $V_{\text{long}}(r)$ is a matrix describing the potential, and its diagonal elements include terms for the mass splittings between states, as in e.g. Hisano:2004ds .
@@
We will assume throughout this section that the potential is spherically symmetric and ignore couplings between states of different $\ell$. Then the Schrödinger equation satisfied by the $\ell$th partial wave of the component of the radial reduced wavefunction in the $n$th state can be written:
@@
where $\hat{V}_{\text{long}}(r)$ is a $N\times N$ matrix that goes to zero as $r\rightarrow\infty$. We will assume throughout that $V_{\text{long}}(r)$ (and hence $\hat{V}_{\text{long}}(r)$) is real and symmetric for all $r$ [^5] The generalization to a complex Hermitian $V_{\text{long}}(r)$ is not difficult; see App. F.
@@ 3.1 The two-body scattering problem for multiple states
As in Sec. 2, this relation assumes distinguishable particles in the initial state, and should be corrected by a prefactor $c_{i}$ which is 1 for distinguishable particles and 2 for identical particles. We will generally work with scattering amplitudes appropriate for distinguishable particles unless noted otherwise (see App. C for a discussion of how to obtain the correctly rescaled matrix elements from QFT for short-distance scattering amplitudes involving identical particles).
@@ 3.4 Extracting the modified $S$-matrix element
Thus we have reduced the problem of computing the full $S$-matrix element to that of computing $S_{0,\ell}$, $\bar{Q}_{\ell}$ and the coefficients in the matrix $R_{\ell}$. $S_{0,\ell}$ and $\bar{Q}_{\ell}$ can be computed using standard methods from the Schrödinger equation with only the long-distance potential. We can also use Eq. 121 to relate these quantities:
@@ 3.5 Characterizing the effects from short-range physics
where $f_{s,\ell}$ is the scattering amplitude describing the short-distance physics (according to our convention, this has no contribution from $V_{\text{long}}(r)$, but $f_{s,\ell}$ may include physical effects that we might usually absorb into the long-range potential). We see that the matching conditions then correspond to choosing $\alpha_{ni}(a)/\sqrt{p_{n}}=\gamma_{ni}$ and $-\beta_{ni}(a)/\sqrt{p_{n}}=\sum_{j\leq N}\gamma_{ji}f_{s,nj}p_{j}$ and obtain
@@
Then we obtain $\tilde{P}f_{s,\ell}\tilde{P}\alpha_{\ell}(a)=-\beta_{\ell}(a)$. Working in matrix notation and decomposing $u_{\ell}(r)$ into its $\tilde{G}_{\ell}(r)$ and $w(r)$ components as Eq. 130, we find that:
@@
×$\times\left(\tilde{P}f_{s,\ell}\tilde{P}\alpha_{w_{\ell}}(a)+\beta_{w_{\ell}}(a)\right).$ (136)
Let $f_{b,\ell}$ be the short-distance scattering amplitude corresponding only to the part of $V_{\text{long}}(r)$ with $r<a$ (i.e. it is a contribution to $f_{s,\ell}$, and may be zero if we choose to set $V_{\text{long}}(r)=0$ for $r<a$). To read off $f_{b,\ell}$, consider a regular solution $(u_{b,\ell})_{ni}$ to the Schrödinger equation with potential $V_{\text{long}}(r)$ for $r<a$ (and matching onto a free-particle solution for $r>a$). Let us denote the $\alpha$ and $\beta$ coefficients associated with this solution by $\alpha_{b,\ell}$, $\beta_{b,\ell}$, and choose the boundary condition at $r=a$ to be $(\alpha_{b,\ell}(a))_{ni}=\sqrt{p_{i}}\delta_{ni}$, or $\alpha_{b,\ell}=\tilde{P}$. This corresponds to sending in a unit-normalized plane wave with component $\delta_{ni}s_{\ell}(p_{n}r)$ in the $n$th component. Then the coefficient of the outgoing $s_{\ell}(p_{n}r)+ic_{\ell}(p_{n}r)$ term at $r=a$ is $-(\beta_{b,\ell}(a))_{ni}/\sqrt{p_{n}}$. For the free-particle propagation at $r>a$, $\alpha_{n}(r)$ and $\beta_{n}(r)$ remain constant, so we can identify the coefficient of this outgoing term with $p_{i}f_{b,ni}$, and $f_{b,\ell}=-\tilde{P}^{-1}\beta_{b,\ell}(a)\tilde{P}^{-2}$, or equivalently $\beta_{b,\ell}(a)=-\tilde{P}f_{b,\ell}\tilde{P}^{2}$.
@@
×$\times\left(\tilde{P}\hat{f}_{s,\ell}\tilde{P}\alpha_{w_{\ell}}(a)\right)$
@@
Now consider the Wronskian between the solutions $\tilde{G}_{\ell}(r)$ and $u_{b,\ell}(r)\tilde{P}^{-1}$ (note this is still a valid solution matrix as the $n$th component of the $i$th solution is $(u_{b,\ell})_{ni}(r)/\sqrt{p_{i}}$, i.e. the $\sqrt{p_{i}}$ rescaling factor is the same constant for all components of any given solution). From Eq. 199 we find that $(\alpha_{b,\ell}(r)\tilde{P}^{-1})^{T}\beta_{\tilde{G}_{\ell}}(r)-(\beta_{b,\ell}(r)\tilde{P}^{-1})^{T}\alpha_{\tilde{G}_{\ell}}(r)$ is a $r$-independent constant. Since $\alpha_{b,\ell}(a)=\tilde{P}$, we have:
@@
×$\times(\alpha_{b,\ell}(0)\tilde{P}^{-1})^{-1}\tilde{P}Q_{\ell}$
@@
$S_{\ell}$ $=S_{0,\ell}\left(1+2i\bar{Q}^{\dagger}_{\ell}\left[\alpha_{b,\ell}(0)\tilde{P}^{-2}\hat{f}_{s,\ell}^{-1}\tilde{P}^{-2}(\alpha_{b,\ell}(0))^{T}\tilde{P}^{2\ell+1}-\alpha_{b,\ell}(0)\tilde{P}^{-1}\alpha_{\tilde{G}_{\ell}}(a)\right.\right.$
@@
Recalling $\bar{Q}_{\ell}=\tilde{P}^{2(\ell+1)}Q_{\ell}P^{-1}$, and defining $\Sigma_{0,\ell}\equiv\tilde{P}^{2}Q_{\ell}P^{-2}$, we can write the $S$-matrix element in the simple form:
@@
We see that our choice of how to define $V_{\text{long}}(r)$ for $r<a$ will affect the $S$-matrix via the $\Sigma_{0,\ell}$ and $S_{0,\ell}$ terms (although the $S_{0,\ell}$ dependence will drop out in terms depending only on $S_{\ell}^{\dagger}S_{\ell}$), and through the coefficient $\kappa_{\ell}$. For example, if we fix $V_{\text{long}}(r)=0$ for $r<a$, then $\alpha_{b,\ell}(0)=\alpha_{b,\ell}(a)=\tilde{P}$, and so we obtain $\kappa_{\ell}^{-1}=\left(\hat{f}_{s,\ell}^{-1}\tilde{P}^{2\ell}-\tilde{P}\alpha_{\tilde{G}_{\ell}}(a)\right)(\tilde{P}^{\dagger})^{-2\ell}$, and $\hat{f}_{s,\ell}=f_{s,\ell}$. So in this convention, the contributions from the short-range part of the non-relativistic potential must be included in $f_{s,\ell}$, which becomes $a$-dependent, and $\Sigma_{0,\ell}$, $S_{0,\ell}$ and $\alpha_{\tilde{G}_{\ell}}(a)$ are also explicitly $a$-dependent. This approach gives the simplest relationship between $f_{s,\ell}$ and $\kappa_{\ell}$, and so can facilitate the calculation of $\kappa_{\ell}$.
@@
As another example, we could choose $V_{\text{long}}(r)$ to be equal to the standard non-relativistic long-range potential for all $r$, including $r<a$. In this case $S_{0,\ell}$ and $\Sigma_{0,\ell}$ are $a$-independent, and $\alpha_{\tilde{G}_{\ell}}(r)$ is independent of $a$ (i.e. $\alpha_{\tilde{G}_{\ell}}(a)$ depends on $a$ only via the evaluation point of the function; the boundary conditions defining the function do not vary as we change $a$). $\alpha_{b,\ell}(0)$ depends on $a$ (via its boundary conditions that are set at $r=a$). $\hat{f}_{s}$ incorporates only short-range ($r<a$) interactions that are not already captured by the long-range potential on its own, and should be $a$-independent at leading order if this physics is indeed short-range; however, higher-order terms involving both $V_{\text{long}}(r<a)$, and the short-range physics not captured in $V_{\text{long}}(r)$, will generally induce an $a$-dependence.
@@
We can also include higher-order real corrections to the potential, including effects that only appear within $r<a$, in $V_{\text{long}}(r)$; in this case $\hat{f}_{s,\ell}$ will no longer include these contributions (they are subtracted by $f_{b,\ell}$), and the factors constructed from the long-range potential ($S_{0,\ell}$, $\Sigma_{0,\ell}$, $\alpha_{\tilde{G}}(a)$, $\alpha_{b,\ell}(0)$) will shift accordingly. The main effect of omitting these higher-order corrections, either in $V_{\text{long}}(r)$ or alternatively in $\hat{f}_{s,\ell}$ (for short-range contributions), is to slightly shift the positions of the resonance peaks (see e.g. Beneke:2019qaa ; Urban:2021cdu ) in parameter space. This shift can drastically change the annihilation cross section at a specified parameter point if that point is sufficiently close to a resonance, but is not expected to change the overall range of possibilities (we demonstrate this point with an example in Sec. 4). Relatedly, we will see that setting $V_{\text{long}}(r)=0$ for $r<a$ changes the pattern of resonance peaks in $\Sigma_{0,\ell}$, which is analogous to the standard Sommerfeld factor; by construction, the additional $a$-dependent terms cancel this shift and restore the correct resonance positions in the full $S$-matrix. This may provide a practical reason to choose a convention in which $V_{\text{long}}(r)$ is $a$-independent, in order to avoid numerical error in the delicate interplay of $a$-dependent factors that controls the correct resonance positions, and consequently in the evaluation of the cross section close to resonance.
@@
We can view $\kappa_{\ell}$ as a corrected short-distance matrix element – it depends on the long-range potential via the $\alpha_{\tilde{G}_{\ell}}(a)$ and $\alpha_{b,\ell}(0)$ terms, but asymptotes to $f_{s,\ell}$ in the limit where $f_{s,\ell}$ and $a$ are both small (or whenever $f_{s,\ell}$ is small if we choose the convention where $V_{\text{long}}(r)=0$ for $r<a$), up to phase factors from the momenta. Similarly, we can regard $\Sigma_{\ell}$ as a corrected Sommerfeld enhancement. As we will discuss in Sec. 3.6.2, the inverse of $\kappa_{\ell}$ is related to $k_{\ell}$ by a simple power of the momentum matrix, so much of our intuition for $k_{\ell}$ from the single-state case can be carried over.

### v3

@@ 2.6.3 Finite range Coulomb potential
In both the $\ell=0$ and $\ell\geq 1$ cases, the divergent behavior of $C_{\ell}^{2}$ as $p\rightarrow 0$ is suppressed by the last factor on the right-hand side of this equation. In Fig. 9, we show the Sommerfeld enhancement as calculated by the conventional formula Eq. 45, the full formula Eq. 42, and the simplified formula Eqs. 99, 100. For the $s$-wave case, the behavior of the full formula is well captured by the simplified formula Eq. 99. For higher partial waves, the simplified formula Eq. 100 describes the behavior of the full formula fairly well, except for the points where $\tilde{z}_{\ell}$ crosses 0 and flips its sign; these behaviors are similar to those for the finite square well.
Figure 9: The ratio between the annihilation cross sections with and without the Sommerfeld effect for the finite-range Coulomb potential, for the $\ell=0$ (left), $\ell=1$ (middle) and $\ell=2$ (right) cases. The black solid curves shows the results computed using the conventional formula in Eq. 45, the red solid curves correspond to the full corrected result given in Eq. 42, and the blue dashed curves correspond to the simplified formula of Eqs. 99, 100. We take $k(p_{0})=(10^{3}+10^{3}i)a_{B}^{-1}$ for the $s$-wave case, $(10^{5}+10^{5}i)a_{B}^{-3}$ for the $p$-wave case, and $(10^{5}+10^{5}i)a_{B}^{-5}$ for the $d$-wave case. We choose $p=5\times 10^{-3}a_{B}^{-1}$ and $a=10^{-10}a_{B}$ in all cases.
@@ 3 The multi-state case for general $\ell$
Assuming the energy associated with the potential to be small compared to the mass, we will generally treat all states that can be excited due to the potential as having approximately equal masses $M_{\chi}$, and treat the mass differences between them as terms in the potential. In this case, $V_{\text{long}}(r)$ is a matrix describing the potential, and its diagonal elements include terms for the mass splittings between states, as in e.g. Hisano:2004ds .
@@
[…]
@@
where $\hat{V}_{\text{long}}(r)$ is a $N\times N$ matrix that goes to zero as $r\rightarrow\infty$. We will assume throughout that $V_{\text{long}}(r)$ (and hence $\hat{V}_{\text{long}}(r)$) is real and symmetric for all rr. [^5] The generalization to a complex Hermitian $V_{\text{long}}(r)$ is not difficult; see App. H.
@@ 3.1 The two-body scattering problem for multiple states
As in Sec. 2, this relation assumes distinguishable particles in the initial state, and should be corrected by a prefactor $c_{i}$ which is 1 for distinguishable particles and 2 for identical particles. We will generally work with scattering amplitudes appropriate for distinguishable particles unless noted otherwise (see App. D for a discussion of how to obtain the correctly rescaled matrix elements from QFT for short-distance scattering amplitudes involving identical particles).
@@ 3.4 Extracting the modified $S$-matrix element
Thus we have reduced the problem of computing the full $S$-matrix element to that of computing $S_{0,\ell}$, $\bar{Q}_{\ell}$, and the coefficients in the matrix $R_{\ell}$. $S_{0,\ell}$ and $\bar{Q}_{\ell}$ can be computed using standard methods from the Schrödinger equation with only the long-distance potential. We can also use Eq. 121 to relate these quantities:
@@ 3.5 Characterizing the effects from short-range physics
where $f_{s,\ell}$ is the scattering amplitude describing the short-distance physics. We see that the matching conditions then correspond to choosing $\alpha_{ni}(a)/\sqrt{p_{n}}=\gamma_{ni}$ and $\beta_{ni}(a)/\sqrt{p_{n}}=\sum_{j\leq N}\gamma_{ji}f_{s,nj}p_{j}$, and obtain
@@
Then we obtain $\tilde{P}f_{s,\ell}\tilde{P}\alpha_{\ell}(a)=-\beta_{\ell}(a)$. Working in matrix notation and decomposing $u_{\ell}(r)$ into its $\tilde{G}_{\ell}(r)$ and $w(r)$ components as in Eq. 130, we find that:
@@
$\times\left(\tilde{P}f_{s,\ell}\tilde{P}\alpha_{w_{\ell}}(a)+\beta_{w_{\ell}}(a)\right).$ (136)
Let $f_{b,\ell}$ be the short-distance scattering amplitude corresponding only to the part of $V_{\text{long}}(r)$ with $r<a$ (i.e. it is a contribution to $f_{s,\ell}$, and may be zero if we choose to set $V_{\text{long}}(r)=0$ for $r<a$). To read off $f_{b,\ell}$, consider a regular solution $(u_{b,\ell})_{ni}$ to the Schrödinger equation with potential $V_{\text{long}}(r)$ for $r<a$ (and matching onto a free-particle solution for $r>a$). Let us denote the $\alpha$ and $\beta$ coefficients associated with this solution by $\alpha_{b,\ell}$, $\beta_{b,\ell}$, and choose the boundary condition at $r=a$ to be $(\alpha_{b,\ell}(a))_{ni}=\sqrt{p_{i}}\delta_{ni}$, or $\alpha_{b,\ell}(a)=\tilde{P}$. This corresponds to sending in a unit-normalized plane wave with component $\delta_{ni}s_{\ell}(p_{n}r)$ in the $n$th component. Then the coefficient of the outgoing $s_{\ell}(p_{n}r)+ic_{\ell}(p_{n}r)$ term at $r=a$ is $(\beta_{b,\ell}(a))_{ni}/\sqrt{p_{n}}$. For the free-particle propagation at $r>a$, $\alpha_{n}(r)$ and $\beta_{n}(r)$ remain constant, so we can identify the coefficient of this outgoing term with $p_{i}f_{b,ni}$, and $f_{b,\ell}=-\tilde{P}^{-1}\beta_{b,\ell}(a)\tilde{P}^{-2}$, or equivalently $\beta_{b,\ell}(a)=-\tilde{P}f_{b,\ell}\tilde{P}^{2}$.
@@
$\times\left(\tilde{P}\hat{f}_{s,\ell}\tilde{P}\alpha_{w_{\ell}}(a)\right)$
@@
Now consider the Wronskian between the wavefunctions $\tilde{G}_{\ell}(r)$ and $u_{b,\ell}(r)\tilde{P}^{-1}$ (note this is still a valid solution matrix as the $n$th component of the $i$th solution is $(u_{b,\ell})_{ni}(r)/\sqrt{p_{i}}$, i.e. the $\sqrt{p_{i}}$ rescaling factor is the same constant for all components of any given solution), which are both solutions to the same Schrödinger equation for $r<a$. From Eq. 212 we find that $(\alpha_{b,\ell}(r)\tilde{P}^{-1})^{T}\beta_{\tilde{G}_{\ell}}(r)-(\beta_{b,\ell}(r)\tilde{P}^{-1})^{T}\alpha_{\tilde{G}_{\ell}}(r)$ is a $r$-independent constant. Since $\alpha_{b,\ell}(a)=\tilde{P}$, we have:
@@
$\times(\alpha_{b,\ell}(0)\tilde{P}^{-1})^{-1}\tilde{P}Q_{\ell}$
@@
[…]
@@
Recalling $\bar{Q}_{\ell}=\tilde{P}^{2(\ell+1)}Q_{\ell}P^{-1}$, let us define $\Sigma_{0,\ell}\equiv\tilde{P}^{2}Q_{\ell}P^{-2}$ (we will see later that this is the standard Sommerfeld matrix). Note that the identity $\bar{Q}_{\ell}^{T}=S_{0,\ell}\bar{Q}_{\ell}^{\dagger}$ now translates to:
$P\Sigma_{0,\ell}^{T}=S_{0,\ell}P\Sigma_{0,\ell}^{\dagger}(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}$ (141)
We can then write the $S$-matrix element in the simple form:
@@
Note that $\alpha_{b,\ell}(0)$ describes the normalization of a regular solution at the origin, so it can be related to the (standard calculation of the) Sommerfeld enhancement if we were to restrict $V_{\text{long}}$ to the region $r<a$ only. Let us write the analogue of $\Sigma_{0,\ell}$ when we set $V_{\text{long}}(r)$ to zero outside $r=a$ as $\Sigma^{a}_{0,\ell}$ (noting however that $\Sigma^{a}_{0,\ell}$ is $N\times N$ rather than $N\times M$). Then using the analogue of Eq. 132 for this case, we can write $\alpha_{b,\ell}(0)=\tilde{P}^{-1}\Sigma^{a}_{0,\ell}\tilde{P}^{2}$, and we can write $\kappa_{\ell}^{-1}$ in the form:
$\kappa_{\ell}^{-1}$ $\equiv\bar{f}_{s,\ell}^{-1}-\bar{Z}_{\ell},\quad\bar{f}_{s,\ell}\equiv(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}(\Sigma^{aT}_{0,\ell})^{-1}\hat{f}_{s,\ell}(\Sigma^{a}_{0,\ell})^{-1},\quad\bar{Z}_{\ell}\equiv\Sigma^{a}_{0,\ell}\tilde{P}\alpha_{\tilde{G}_{\ell}}(a)(\tilde{P}^{\dagger})^{-2\ell}.$ (143)
In general, we see that $\bar{f}_{s,\ell}$ encodes the short-range physics, potentially including effects from evolution under $V_{\text{long}}(r)$ at $r<a$. In contrast, $\bar{Z}_{\ell}$ depends solely on $V_{\text{long}}(r)$.
We show in App. E that if we were to model the physics at $r<a$ as a contact interaction (potentially including both elastic and inelastic interactions) combined with whatever physics is already captured in $V_{\text{long}}(r)$ for $r<a$, and work to lowest order in the contact interaction, then $\bar{f}_{s,\ell}$ would be fixed by the scattering amplitude associated with the contact interaction, and hence would be $a$-independent. Roughly speaking, the factor of $(\Sigma^{a}_{0,\ell})^{-1}$ and its transpose take the full scattering amplitude and factor out the Sommerfeld factors arising from the already-modeled $V_{\text{long}}(r)$ at $r<a$, leaving only the raw contact-interaction contribution. In a similar way, to obtain $\hat{f}_{s}$ we already subtracted off the scattering amplitude arising solely from interactions involving $V_{\text{long}}(r)$ at $r<a$. Thus we may understand $\bar{f}_{s,\ell}$ as capturing the $r<a$ physics but stripping off the leading additive and multiplicative terms associated with $V_{\text{long}}(r)$ at $r<a$.
We see that our choice of how to define $V_{\text{long}}(r)$ for $r<a$ will in general affect the $S$-matrix via the $\Sigma_{0,\ell}$, $S_{0,\ell}$, $\bar{f}_{s,\ell}$ and $\bar{Z}_{\ell}$ terms (although the $S_{0,\ell}$ dependence will drop out in terms depending only on $S_{\ell}^{\dagger}S_{\ell}$). There are two particularly simple conventions one might choose.
One choice would be to fix $V_{\text{long}}(r)=0$ for $r<a$, so that $\bar{f}_{s,\ell}$ explicitly includes all contributions from the short-range part of the non-relativistic potential, and so will be $a$-dependent. In this case, there is no potential for $r<a$, so $\Sigma^{a}_{0,\ell}=1$, $f_{b,\ell}=0$, and consequently $\bar{f}_{s,\ell}=(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}f_{s,\ell}$. The $\Sigma_{0,\ell}$, $S_{0,\ell}$ and $\bar{Z}_{\ell}$ factors are also explicitly $a$-dependent. This convention choice corresponds to the simplest definition of what it means to separate “short-range” ($r<a$) and “long-range” ($r>a$) physics. We will find it helpful for proving a number of results in the following sections, as it gives the simplest expressions for $\bar{f}_{s,\ell}$ and $\bar{Z}_{\ell}$.
@@
An alternative choice is to set $V_{\text{long}}(r)$ to be equal to the non-relativistic long-range potential for all $r$, including $r<a$, such that the definition of $V_{\text{long}}(r)$ does not change when we vary $a$. In this case $S_{0,\ell}$ and $\Sigma_{0,\ell}$ are $a$-independent, and $\alpha_{\tilde{G}_{\ell}}(r)$ is independent of $a$ as a function (i.e. $\alpha_{\tilde{G}_{\ell}}(a)$ depends on $a$ only via the evaluation point of the function; the boundary conditions defining the function do not vary as we change $a$). In this convention choice, if the short-range physics not captured in $V_{\text{long}}(r)$ were modeled as a contact interaction, then (as discussed above) $\bar{f}_{s,\ell}$ would be determined solely by the amplitude for the contact interaction (at least at leading order), and hence would be $a$-independent. However, higher-order terms involving both $V_{\text{long}}(r<a)$, and the short-range physics not captured in $V_{\text{long}}(r)$, could induce an $a$-dependence.
To the degree that we model the additional (non-potential) short-range physics as a contact interaction computed through perturbative quantum field theory (i.e. by taking the low-momentum limit), we can match the resulting amplitude directly onto $\bar{f}_{s,\ell}$ in the 2nd convention choice (with a subtlety due to the $(\tilde{P}^{\dagger}\tilde{P}^{-1})^{2\ell}$ term in Eq. 143 that is discussed in App. D). If we switch to the 1st convention choice, $\bar{f}_{s,\ell}$ should also include additional terms from scattering in the potential at $r<a$ and the short-range Sommerfeld factors, as described in App. E. These factors can be computed perturbatively in QFT, or by solving the wavefunction evolution in quantum mechanics for the truncated potential.
In either convention, there is also a choice as to the order at which one computes the non-relativistic potential $V_{\text{long}}(r)$. Provided the higher-order corrections to $V_{\text{long}}(r)$ are real and symmetric (or Hermitian, if the more general results of App. H are used), there is no difficulty in including them in $V_{\text{long}}(r)$; alternatively, if they have suppressed support outside $r=a$, they can be folded into $f_{s,\ell}$. If these higher-order corrections are omitted entirely, the main effect is to slightly shift the positions of the resonance peaks (see e.g. Beneke:2019qaa ; Urban:2021cdu ) in parameter space. This shift can drastically change the annihilation cross section at a specified parameter point if that point is sufficiently close to a resonance, but is not expected to change the overall range of possibilities (we demonstrate this point with an example in Sec. 4).
@@
Relatedly, we will see that setting $V_{\text{long}}(r)=0$ for $r<a$ changes the pattern of resonance peaks in $\Sigma_{0,\ell}$, which is analogous to the standard Sommerfeld factor; by construction, the additional $a$-dependent terms cancel this shift and restore the correct resonance positions in the full $S$-matrix. This provides one practical reason to choose a convention in which $V_{\text{long}}(r)$ is $a$-independent, in order to avoid numerical error in the delicate interplay of $a$-dependent factors that controls the correct resonance positions, and consequently in the evaluation of the cross section close to resonance.
@@
We can view $\kappa_{\ell}$ as a corrected short-distance scattering amplitude – it depends on the long-range potential via the $\bar{Z}_{\ell}$ term, but asymptotes to $\bar{f}_{s,\ell}$ in the limit where $\bar{f}_{s,\ell}$ is small. Similarly, we can regard $\Sigma_{\ell}$ as a corrected Sommerfeld enhancement. As we will discuss in Sec. 3.6.2, the inverse of $\kappa_{\ell}$ is related to $k_{\ell}$ by a simple power of the momentum matrix, so much of our intuition for $k_{\ell}$ from the single-state case can be carried over.